OpenStudy (anonymous):

Solve the equation using the quadratic formula. 2x2 = -12x - 3

OpenStudy (anonymous):

You need to use the quadratic formula and set it up in standard form. -------------------------------------------------------------- Hey I'm doing a survey on music! Can you help me out by answering a couple of questions? Thank you ! http://openstudy.com/study#/updates/503732e9e4b08ecf834fa73a

hero (hero):

Hint: Let a = 2 b = 12 c = 3 Then use the quadratic formula: \(\huge x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\)

OpenStudy (anonymous):

2*x2=12*-3 Moved these terms to the left 12*x,3 It becomes 2*x2+12*x+3=0